### v1

@@ 4.5 Beyond the path-star setting
(Clever Hans cheating due to teacher-forcing) Although there is a true mechanism that can recover each $r_{i}$ from the original prefix $\bm{p}$, there can be multiple other mechanisms that can recover a token $r_{i}$ from the teacher-forced prefix $(\bm{p},\bm{r}_{<i})$. These mechanisms can be simpler to learn thus disincentivizing the model from learning the true mechanism.
@@
(Indecipherable token due to lost supervision) After the Clever Hans cheat is perfected during training, the model is deprived of a part of the supervision (especially, $r_{i}$ for larger $i$). This makes it harder and potentially even intractable for the model to learn the true mechanism from the remaining tokens alone.
As we demonstrate in the next section, the above failures can cause the model to fail on the very distribution it was trained on. This breakdown of planning abilities emerges right from training, and is orthogonal to the Snowballing Failure that is primarily an inference-time issue (See §A).
While the path-star problem provides a concrete and empirically verifiable setting of this failure (where it is easy to reason and test the mechanisms), it can help us speculate how such failures could occur in more complex and nebulous tasks. Intuitively we expect this failure to occur when there are right-to-left dependencies: a token that appears later must be planned before an earlier-appearing token. We provide an example below.
Story-writing. Consider training with teacher-forcing on novels. Imagine that the (sub)plots take the form of a conflict, followed by a backstory, followed by a resolution of the conflict, utilizing the backstory (illustrated in detail in §B). Crucially, although the story explicitly reads as $v_{\tt conflict},v_{\tt backstory},v_{\tt resolution}$, the model must learn to decide on $v_{\tt backstory}$ before all else.
However, we hypothesize that a teacher-forced model trained would fail to learn this story-writing plan. First, the teacher-forced model would utilize the Clever Hans cheat: it would learn the deductive skills required to fit the last segment $v_{\tt resolution}$ using the preceding segments revealed by the teacher in the input, $v_{\tt conflict},v_{\tt backstory}$. With the supervision from $v_{\tt resolution}$ lost, the model has no explicit indication of how $v_{\tt conflict}$ and $v_{\tt backstory}$ depend on each other. These would become Indecipherable Tokens. We hypothesize that the resulting model would learn to generate uninteresting stories, interjecting arbitrary conflicts and backstories on a whim, subsequently forcing contrived resolutions upon them. While this hypothesis is not straightforward to empirically test for, we provide a more detailed conceptual illustration in §B.
@@ 5 Experimental Verification
In this section, we demonstrate our hypothesized failure modes in practice on the graph path-finding task. We perform our experiments in both Transformers and Mamba to demonstrate that these failures are general to teacher-forced models. We begin by establishing that our teacher-forced models fit the training data but fail in-distribution. Next, we design metrics to quantify the extent to which the two hypothesized mechanisms occur (Failures 2a2b). Finally, we design alternative objectives to intervene and remove each of the two failure modes, to test whether the performance improves. We report additional experiments in §D.1 quantifying the Snowballing Failure 1. We describe our experimental setting more precisely below.
@@
We denote by ${G}_{d,l}(N)$ for $d,l,N\in\mathbb{N}$, a path-star graph consisting of a center node ${v}_{\tt{start}}$ with degree $d\in\mathbb{N}$, meaning there are $d$ different paths emerging from the center node, each consisting of $l-1$ nodes (excluding the start node). Node values are uniformly sampled from ($(\{0,\dots,N-1\})$ where $N$ can be larger than the actual number of nodes in the path-star graph. In every graph, we use the center node as the starting node ${v}_{\tt{start}}$ and then pick as ${v}_{\tt{goal}}$, the last node of one of the paths chosen uniformly at random. The order of the edges in the adjacency list is randomized. We describe the tokenization in §E.1.
@@
We evaluate models from two architectural families to highlight that the failures are not tied to a particular architecture but stem from the next-token prediction objective. For Transformers, we use from-scratch GPT-Mini, and pretrained GPT-2 large (Radford et al., 2019). For recurrent models, we use from-scratch Mamba (Gu & Dao, 2023). We optimize using AdamW (Loshchilov & Hutter, 2019) until perfect training accuracy. To rule out grokking behaviour (Power et al., 2022), we trained the cheaper models for as long as $500$ epochs. More details are in §E.2.
@@
For a given distribution, we evaluate all our teacher-forced models by autoregressively generating solutions, and comparing that solution with the true one for an exact-match:
$\texttt{Acc}_{\texttt{ag}}(\texttt{LM}_{\theta}):=\mathbb{P}(\hat{\bm{r}}=\bm{r})\text{,}$ $\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}.$ (4)
We report $\texttt{Acc}_{\texttt{ag}}(\texttt{LM}_{\theta})$ for path-star graphs of varying topologies in Fig. 3 and Table 2. We observe that all models (even when pre-trained) struggle to learn the task accurately. The accuracy values are precisely limited to the value achievable if the model uniformly guesses a path starting from ${v}_{\tt{start}}$ i.e., $\approx\frac{1}{d}$, thus establishing complete in-distribution failure. This is so even when trained to fit sample sizes up to $200k$ to $100\%$ accuracy, and despite the fact that the training and test graphs have identical topology. Next, we quantitatively demonstrate how this stark failure arises from our two hypothesized mechanisms (Failure 2a2b).
@@
We had hypothesized that the teacher-forcing model would use cheating to fit the training tokens (the ones that follow $\bm{r}_{1}$ in each instance). Specifically, to predict node $v_{i}$ in the true path, the model can exploit the ground truth node $v_{i-1}$ that is revealed as input. Rather than learning to plan, the model would simply predict the node that is outwardly adjacent to $v_{i-1}$. To quantify whether this behavior emerges, we “teacher-force” the model with a uniform random neigbhor $v_{1}^{\prime}$ of ${v}_{\tt{start}}$. We then look for whether the model indiscriminately applies the learned Clever Hans cheat here: does the model simply follow along the path that emanates from the neigbhor $v_{1}^{\prime}$, not necessarily ending in ${v}_{\tt{goal}}$?
$\hskip 5.69054pt\text{where}\hskip 8.53581pt\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}_{1<}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot;\bm{p},{v}_{\tt{start}},v_{1}^{\prime})$
Empirically, we find that $\texttt{Acc}_{\texttt{cheat}}(\texttt{LM}_{\theta})$ on a held-out test set is $\approx 100\%$ almost across the board (except for graphs with very high degree where training is challenging). The exact values are in §5, Table 1. This establishes that to fit the training data, the teacher-forced model has exploited the Clever Hans cheat.
@@
Recall that the Clever Hans cheat only applies to all but the first node $v_{1}$ after ${v}_{\tt{start}}$ lying on the path. After the Clever Hans cheat fits the rest of the path during training, we hypothesized that node $v_{1}$ may become impossible to learn since the model is deprived of all information about the subsequent targets. To quantify this behavior, we evaluate how well the model is able to predict the difficult first node, $v_{1}$:
$\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})=\mathbb{P}\left(\hat{r}_{1}=r_{1}\right),$ $\bm{p},\bm{r}\sim\mathcal{D},\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot;\bm{p}).$ (6)
which we estimate using the held-out test set. As shown in Fig. 4 the model achieves a low $\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})$, approximately $1/d$. Thus, the model indeed fails to identify that $v_{1}$ is the one on the path to ${v}_{\tt{goal}}$. It instead resorts to emitting one of the $d$ neighbors of ${v}_{\tt{start}}$ at random.
@@
We denote a model trained in this fashion by $\texttt{LM}_{\theta}^{\$}$ and perform inference simply by conditioning on \$currency-dollar\\$\$ tokens i.e.
$\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt\$}}}}}{{\sim}}\texttt{LM}_{\theta}^{\$}(\cdot;\bm{p})\text{ where }\hat{r}_{i}\sim\texttt{LM}_{\theta}^{\$}(\cdot;\bm{p},\bm{r}^{\$}_{<i}).$
and accordingly evaluate the accuracy:
$\texttt{Acc}_{\texttt{\$}}(\texttt{LM}_{\theta}^{\$})=\mathbb{P}\left(\hat{\bm{r}}=\bm{r}\right)$ $\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt\$}}}}}{{\sim}}\texttt{LM}_{\theta}^{\$}(\cdot;\bm{p}).$ (8)
This training and inference setup was proposed in Monea et al. (2023) for the orthogonal purpose of improving the computational efficiency of inference. Our goal however is to evaluate whether forcing the model to lookahead can prevent the Clever Hans cheat from being picked up, and thereby allow the model to generalize successfully. We report the accuracy of these teacherless models in Fig. 3 and Table 3. Unfortunately, in most cases, the teacherless objective is too hard for the models to even fit the training data, likely because there is no simple cheat to employ here. However, surprisingly, on some of the easier graphs, the models not only fit the training data, but generalize well to test data. This positive result (even if in limited settings) verifies two hypotheses. First, the Clever Hans cheat is indeed a cause of failure in the original teacher-forced model. Secondly, and remarkably, with the cheat gone, these models are able to fit the first node which had once been indecipherable under teacher-forcing. This verifies our hypothesis that the Clever Hans cheat absorbs away supervision that is critical to learn the first token. Shortly at the end of this section, we provide more intuition about how exactly teacherless models are able to solve this task with the Clever Hans cheat out of their way.
@@
Notationally, we let $\texttt{LM}_{\theta}^{\texttt{rev}}$ be the model trained to maximize $\mathcal{J}_{\texttt{next-token}}$ with the targets (and the teacher-forced inputs) set to $\bm{r}^{\texttt{rev}}=r_{L_{\texttt{resp}}},\ldots r_{1}$, the reversal of $\bm{r}$. We then measure the autoregressive accuracy by comparing against $\bm{r}^{\texttt{rev}}$:
$\texttt{Acc}_{\texttt{rev}}(\texttt{LM}_{\theta}^{\texttt{rev}})=\mathbb{P}\left(\hat{\bm{r}}=\bm{r}^{\texttt{rev}}\right),\;\bm{p},\bm{r}\sim\mathcal{D},\hat{\bm{r}}\stackrel{{\scriptstyle\mathclap{\mbox{{\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}^{\texttt{rev}}(\cdot;\bm{p})$ (9)
@@
We conclude by emphasizing that the success of the reversed training (and also, the occasional success of teacherless training) make the in-distribution failure of teacher-forcing particularly surprising. Recall that, when viewed left-to-right, our problem requires complex planning — evaluating multiple paths and selecting the right one — but when viewed right-to-left, the problem is straightforward. The experiments on the reversed formulation confirm that the right-to-left solution is not only expressible by our architectures, but also learnable via gradient descent. Evidently, the left-to-right teacher-forced model is unable to view the problem any differently and falls into the traps outlined in §4.
Intuition for teacherless training. We hypothesize that even teacherless training allows the model to implicitly learn the right-to-left view. Concretely, the teacherless model cannot use the trivial Clever Hans cheat to fit the data, since the ground truth prefixes are not available during training. Nor is it explicitly prescribed to fit the target right-to-left. Instead the model is tasked with using only the graph description in $\bm{p}$ to fit all the target nodes $\bm{r}$ (implicitly requiring a lookahead beyond just the next token). In this paradigm, the model would first fit the target token that is simplest to deduce using only information available in the prefix $\bm{p}$: this is the penultimate vertex $r_{l-1}$ which is the unique token that precedes the goal (and can be discovered using a simple scan of the prefix). Once the model figures this out, the model can similarly work backwards to fit each node $r_{i-1}$ using the previously-fit $r_{i}$. (We note that this solution is still a fairly difficult one to implement in the teacherless setting, which surprisingly some models nevertheless learn — we describe this in §C.)
Our hypothesis is borne out in Fig. 4 where we see that the teacherless model automatically learns right-to-left: the later tokens achieve higher accuracy earlier. Thus, the teacherless objective provides one possible way for future work to build alternatives to next-token prediction that force the models to look ahead, without falling into the next-token traps of the Clever Hans cheat or the Indecipherable Token failure.
@@ 6 Related Work
We provide an elaborate survey of the various arguments that have been made in favor of or against next-token prediction. We hope this can help consolidate a debate that has been scattered over the literature. Part of the survey is deferred to §F.
@@
Arguments in support of next-token prediction. Shannon (1948; 1951); Alabdulmohsin et al. (2024) demonstrate that language has enough redundancy to be conducive for next-token prediction. Empirically, Shlegeris et al. (2022) find that modern language models are surprisingly better than humans at next-token prediction on the text dataset, OpenWebText (Gokaslan & Cohen, 2019). But this does not preclude the possibility that next-token predictors may still be poor at planning. Furthermore, the above result may be confounded by the ability of language models to store more general knowledge than humans.
@@
On the theoretical side, Merrill & Sabharwal (2023b); Feng et al. (2023) show that autoregressive Transformers that generate chains of thought have a markedly larger expressive power. Most relevant to us is the positive learnability results of Malach (2023); Wies et al. (2023) which argue that complex multi-hop tasks that are otherwise unlearnable, become learnable via next-token prediction when there is a preceding chain-of-thought supervision for each hop. Our negative result does not contradict this. In our path-finding problem, learning the first token requires an implicit chain of thought (the reversed path) that we do not provide before the first token.
Arguments against next-token prediction. The most well-formulated criticism of next-token prediction is what we term as the snowballing error, scattered in literature, both in recent work Dziri et al. (2023); LeCun (2024) and much earlier in Kääriäinen (2006); Ross & Bagnell (2010). The earlier works capture an additional notion of snowballing, wherein, once an erroneuous sub-optimal action is committed, the model is more likely to commit more sub-optimal actions since it has wandered into territories that it was not trained on. Implicitly, the error here is not evaluated as an exact match of the response (i.e., $\bm{r}\neq\hat{\bm{r}}$) but as a cumulative notion of error over all steps (e.g., $\sum\mathbf{1}[r_{i}\neq\hat{r}_{i}]$). In this setting, there is an additional cause of failure called exposure bias: the teacher-forced model has only been trained on correct trajectories, and has not learned how to recover from poor trajectories. Nevertheless, all existing criticisms assumes that teacher-forcing has learned an accurate next-token predictor in the first place, which is the very assumption our counterexample challenges. Indeed, in our setting the error happens “instantaneously” at the beginning of inference.
@@
Our main counterexample can be seen as a way of formalizing an emerging, informal intuition that is often worded as “autoregressive next-token predictors are ill-suited for planning tasks”. Indeed, Momennejad et al. (2023); Valmeekam et al. (2023) report failures on several planning tasks framed as word problems (including path-finding in Momennejad et al. (2023)) and Bubeck et al. (2023) on various arithmetic, summarization and poem/story generation problems (which they speculate is a limit of autoregressivity). McCoy et al. (2023) argue that, for such tasks, the performance of the model must greatly depend on its frequency during pretraining. However, we demonstrate that even when trained on many samples from a distribution, the next-token predictor can fail on the very distribution.
A closely-related criticism (Bubeck et al., 2023; Dawid & LeCun, 2023; LeCun, 2024; Du et al., 2023a) is that to model human thinking, we need to model two types of thinking as outlined in Kahneman (2011): a fast (System 1) thinking process that is also guided by a slower (Sytem 2) thinking process. Theoretically, Lin et al. (2021) show that there are formal languages for which expressing some next-tokens may require super-polynomial time or parameter count during inference. Du et al. (2023a) informally note that some next tokens can be hard to learn as they require a global understanding of what will be uttered in the future.  [^4] The Indecipherable Token failure is related to the “locally unlearnable” hypothesis of Du et al. (2023a), but is not the same. The (first) Indecipherable Token in the path-star problem is locally learnable by the teacherless model, even without access to the full target sequence (which is only presented after this token). Thus, the inability of the teacher-forced model to learn the Indeciperhable Token is crucially related to the fact that the supervisory targets that come after this token are lost to the Clever Hans cheat.
@@
Our work extends and clarifies this discourse by introducing the Clever Hans cheat and the Indecipherable Token failure. Next, we empirically report our failure modes in both the Transformer (Vaswani et al., 2017) and the Mamba structured state space model (Gu & Dao, 2023). This establishes that what we witness is indeed a failure of next-token prediction (and not of the Transformer architecture as some existing criticisms are framed). Importantly, existing literature pins these failures broadly on the next-token prediction paradigm and interchangeably, on the inability of the autoregressive architecture to backtrack. We emphasize the need to differentiate between the two types of next-token prediction (teacher-forcing and autoregressive inference) as they lead to distinct planning-related failures and require distinct solutions.
Going beyond next-token prediction. Inference-time techniques like chain-of-thought and its variants (Wei et al., 2022; Yao et al., 2023a; Besta et al., 2023; Yao et al., 2023b) or those that elicit feedback from the model (Madaan et al., 2023; Huang et al., 2022; Shinn et al., 2023) can be thought of as going beyond conventional form of inference by allowing the model to think more before producing its final answer. However, the backbone in these models are still trained by standard teacher-forcing. While other techniques (Burtsev et al., 2020; Xue et al., 2023; Goyal et al., 2023) train the model to explicitly think more, even these boil down to next-token prediction during training.

### v2

@@ 4.5 Beyond the Path-Star Setting
(Clever Hans cheating due to teacher-forcing) Although there is a true mechanism that recovers each $r_{i}$ from the original prefix $\bm{p}$, there may be multiple other mechanisms that can recover each token $r_{i}$ from the teacher-forced prefix $(\bm{p},\bm{r}_{<i})$. These mechanisms may be simpler thus disincentivizing the model from learning the true mechanism.
@@
(Indecipherable token due to lost supervision) After the Clever Hans cheat is perfected during training, the model is deprived of a part of the supervision (especially, $r_{i}$ for larger $i$). This makes it harder and potentially even intractable to learn the true mechanism from the remaining tokens alone.
As we demonstrate in the next section, the above failures can cause the model to fail on the very distribution it was trained on. This breakdown of planning abilities emerges right from training, and is orthogonal to the Snowballing Failure that is primarily an inference-time issue (See §B).
While the path-star problem provides a concrete, verifiable setting of this failure, it can also help us speculate how such failures could occur in more complex and nebulous tasks. More generally, we expect this failure to occur when there are right-to-left dependencies i.e., a later-appearing token must be planned before an earlier-appearing token. We provide an example below.
Story-writing. Imagine training on novels that take the form of a conflict, followed by a backstory, followed by a resolution of the conflict, utilizing the backstory. Although the story explicitly reads as $v_{\tt conflict},v_{\tt backstory},v_{\tt resolution}$, implicitly, one must learn to decide on $v_{\tt backstory}$ before all else. We however conjecture that the teacher-forced model would suffer the Clever Hans cheat wherein it would first learn to fit $v_{\tt resolution}$ using simple deductive skills. With this crucial part of the story lost as supervision, the model can no longer decipher how the remaining pieces relate i.e., how $v_{\tt backstory}$ must be planned in advance of $v_{\tt conflict}$. We conjecture that the resulting model would learn to generate uninteresting stories, interjecting arbitrary conflicts and backstories on a whim, subsequently forcing contrived resolutions upon them. While this hypothesis is not straightforward to empirically test for, we provide a more detailed conceptual illustration in §C.
@@ 5 Experimental Verification
In this section, we demonstrate our hypothesized failure modes on the graph path-finding task. We show this in both Transformers and Mamba to demonstrate that these failures are general to teacher-forced models. First, we establish that our teacher-forced models fit the training data but fail in-distribution. Next, we design metrics to quantify the extent to which the two hypothesized mechanisms (Failures 2a2b) occur. Finally, we design alternative objectives to intervene and remove each of the two failure modes, to test whether the performance improves. We report additional experiments in §F.4 for an arithmetic task, and in §F.1 quantifying the Snowballing Failure 1. We describe our experimental setting more precisely below.
@@
We denote by ${G}_{d,l}(N)$ for $d,l,N\in\mathbb{N}$, a path-star graph consisting of a center node ${v}_{\tt{start}}$ with degree $d\in\mathbb{N}$, meaning there are $d$ different paths emerging from the center node, each consisting of $l-1$ nodes (excluding the start node). Node values are uniformly sampled from ($(\{0,\dots,N-1\})$ where $N$ can be larger than the actual number of nodes in the path-star graph. In every graph, we use the center node as the starting node ${v}_{\tt{start}}$ and then pick as ${v}_{\tt{goal}}$, the last node of one of the paths chosen uniformly at random. The order of the edges in the adjacency list is randomized. We describe the tokenization in §G.1.
@@
We evaluate models from two architectural families to highlight that the failures are not tied to a particular architecture but stem from the next-token prediction objective. For Transformers, we use from-scratch GPT-Mini, and pretrained GPT-2 large (Radford et al., 2019). For recurrent models, we use from-scratch Mamba (Gu & Dao, 2023). We optimize using AdamW (Loshchilov & Hutter, 2019) until perfect training accuracy. To rule out grokking behaviour (Power et al., 2022), we trained the cheaper models for as long as $500$ epochs. More details are in §G.2.
@@
For a given distribution, we evaluate all our teacher-forced models by autoregressively generating solutions, and comparing that solution with the true one for an exact-match. We denote this accuracy as $\texttt{Acc}_{\texttt{ag}}(\texttt{LM}_{\theta})$ (see Eq 5) and report it for path-star graphs of varying topologies in Fig. 3 and Table 2. As observed, all models (even when pre-trained) struggle to learn the task accurately. The accuracies are precisely limited to the value when uniformly guessing a path from ${v}_{\tt{start}}$ i.e., $\approx\frac{1}{d}$, thus establishing complete in-distribution failure. This is so even when trained to fit sample sizes up to $200k$ to $100\%$ accuracy, and despite the fact that the training and test graphs have identical topology. Next, we quantitatively demonstrate how this stark failure arises from our two hypothesized mechanisms (Failure 2a2b).
@@
We had hypothesized that the teacher-forced model would cheat to fit the training tokens (the ones that follow $r_{1}$ in each instance). Specifically, to predict node $v_{i}$ in the true path, the model can exploit the ground truth node $v_{i-1}$ that is revealed as input. Rather than learning to plan, the model would simply predict the node that is outwardly adjacent to $v_{i-1}$. To quantify whether this behavior emerges, we “teacher-force” the model with a uniform random neigbhor $v_{1}^{\prime}$ of ${v}_{\tt{start}}$. We then test whether the model indiscriminately applies the learned Clever Hans cheat here: does the model religiously follow the path that emanates from the neigbhor $v_{1}^{\prime}$, not necessarily ending in ${v}_{\tt{goal}}$? We measure the exact match of this path on a held-out set as $\texttt{Acc}_{\texttt{cheat}}(\texttt{LM}_{\theta})$ ( Eq 6).
Empirically, in §5, Table 1, we find $\texttt{Acc}_{\texttt{cheat}}(\texttt{LM}_{\theta})\approx 100\%$ almost across the board (except for high-degree graphs where training is challenging). This establishes that to fit the training data, the teacher-forced model has exploited the Clever Hans cheat.
@@
Recall that the Clever Hans cheat only applies to all but the first node $v_{1}$ after ${v}_{\tt{start}}$ lying on the path. After the Clever Hans cheat fits the rest of the path during training, we hypothesized that node $v_{1}$ may become impossible to learn since the model is deprived of all information about the subsequent targets. To quantify this behavior, we evaluate how well the model is able to predict the difficult first node, $v_{1}$. We measure this on the held-out set and denote this as $\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})$ (see Eq 7). As shown in Fig. 4 the model achieves a low $\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})$, approximately $1/d$. Thus, the model indeed fails to identify that $v_{1}$ is the one on the path to ${v}_{\tt{goal}}$ and instead randomly emitting one of the $d$ neighbors of ${v}_{\tt{start}}$.
@@
We denote a model trained this way by $\texttt{LM}_{\theta}^{\$}$ and perform inference simply by conditioning on \$currency-dollar\\$\$ tokens i.e., to extract $\hat{r}_{i}$, we feed the uninformative prefix $\bm{r}^{\$}_{<i}$ as input, rather than autoregressively feeding the output $\hat{\bm{r}}_{<i}$ as input. We denote the resulting accuracy by $\texttt{Acc}_{\texttt{\$}}(\texttt{LM}_{\theta}^{\$})$ (see Eq 9). This training and inference setup was proposed in Monea et al. (2023) for the orthogonal goal of improving the computational efficiency of inference. Our goal however is to evaluate whether forcing the model to lookahead can dodge the Clever Hans cheat, thereby allowing the correct mechanism to be learned.
We report the accuracy of these teacherless models in Fig. 3 and Table 3. Unfortunately, in most cases, the teacherless objective is too hard for the models to even fit the training data, likely because there is no simple cheat to employ here. However, surprisingly, on some of the easier graphs, the models not only fit the training data, but generalize well to test data. This positive result (even if in limited settings) verifies two hypotheses. First, the Clever Hans cheat is indeed caused failure in the original teacher-forced model. Secondly, and remarkably, with the cheat gone, these models are able to fit the first node which had once been indecipherable under teacher-forcing. This verifies our hypothesis that the Clever Hans cheat absorbs away supervision that is critical to learn the first token. At the end of this section, we provide more intuition for how the absence of Clever Hans cheat, allows the teacherless models to solve this task.
@@
The success of the reversed training (and of teacherless training) make the in-distribution failure of teacher-forcing particularly surprising. When viewed left-to-right, our problem requires complex planning — evaluating multiple paths and selecting the right one — but when viewed right-to-left, the problem is straightforward; the experiments on the reversed formulation confirm that the right-to-left solution is not only expressible by our architectures, but also learnable via gradient descent. Evidently, the left-to-right teacher-forced model is unable to view the problem any differently and falls into the traps outlined in §4.
Intuition for teacherless training. We hypothesize that even teacherless training allows the model to implicitly learn the right-to-left view. Concretely, the teacherless model cannot use the trivial Clever Hans cheat to fit the data, since the ground truth prefixes are not available during training. Nor is it explicitly prescribed to fit the target right-to-left. Instead the model is tasked with using only the graph description in $\bm{p}$ to fit all the target nodes $\bm{r}$ (implicitly requiring a lookahead beyond just the next token). Our key intuition is that, in this paradigm, the model would first fit the target token that is simplest to deduce using only information available in the prefix $\bm{p}$: this is the penultimate vertex $r_{l-1}$ which is the unique token that precedes the goal (and can be discovered using a simple scan of the prefix). Once the model figures this out, the model can similarly work backwards to fit each node $r_{i-1}$ using the previously-fit $r_{i}$. (Also see Remark 2)
Our hypothesis is borne out in Fig. 4 where we see that the later tokens achieve higher accuracy earlier, implying that the teacherless model voluntarily learns right-to-left. Thus, the teacherless objective provides an alternative training paradigm that forces models to look ahead, without falling into the various short-sighted pitfalls of next-token prediction, discussed in §4.
@@ 6 Related Work
We consolidate the arguments surrounding next-token prediction that has been fragmented over various lines of works. Part of our elaborate survey is deferred to §H.
@@
Arguments in support of next-token prediction. Shannon (1948, 1951); Alabdulmohsin et al. (2024) demonstrate that language has enough redundancy to be conducive for next-token prediction. Empirically, Shlegeris et al. (2022) find that modern language models are surprisingly better than humans at next-token prediction on the text dataset, OpenWebText (Gokaslan & Cohen, 2019). But this does not preclude the possibility that next-token predictors may still be poor at planning. Furthermore, the above result may be confounded by the ability of language models to store more general knowledge than humans.
@@
On the theoretical side, Merrill & Sabharwal (2024); Feng et al. (2023) show that autoregressive Transformers that generate chains of thought have a larger expressive power. Most relevant to us is the positive learnability results of Malach (2023); Wies et al. (2023) which argue that complex multi-hop tasks that are otherwise unlearnable, become learnable via next-token prediction when there is a preceding chain-of-thought supervision for each hop. Our negative result does not contradict this. In our problem, learning the first token requires an implicit chain of thought (the reversed path) that we do not provide before the first token.
Arguments against next-token prediction. The most well-formulated criticism is the snowballing failure mode, which appears scattered in various forms in literature Dziri et al. (2024); LeCun (2024); Kääriäinen (2006); Ross & Bagnell (2010). As explained earlier, this is orthogonal to our failure; indeed, in our setting the error happens “instantaneously” at the beginning, rather than snowball over time.
@@
Our main counterexample can be seen as formalizing an emerging, informal intuition, often worded as “autoregressive next-token predictors are ill-suited for planning tasks”. Indeed, Momennejad et al. (2023); Valmeekam et al. (2023a, b, c) report failures on several planning tasks framed as word problems (including path-finding in Momennejad et al. (2023)) and Bubeck et al. (2023) on various arithmetic, summarization and poem/story generation tasks. McCoy et al. (2023) argue that, for such tasks, the performance of the model must greatly depend on its frequency during pretraining. However, we show that even when trained on many samples from a distribution, the next-token predictor can fail on the very distribution.
@@
Our work extends and clarifies this discourse by introducing the Clever Hans cheat and the Indecipherable Token failure. Next, we empirically report our failure modes in both the Transformer (Vaswani et al., 2017) and the Mamba structured state space model (Gu & Dao, 2023). Thus, what we witness is indeed a failure of next-token prediction (and not of the Transformer architecture as some existing criticisms are framed). Importantly, existing literature pins these failures broadly on the next-token prediction paradigm and interchangeably, on the inability of the autoregressive architecture to backtrack. We emphasize the need to differentiate between the two types of next-token prediction (teacher-forcing and autoregressive inference) as they lead to distinct planning-related failures and require distinct solutions.